# A classification result for eternal mean convex flows of finite total curvature type

### 18 Mar 2024 v1

@@ 3. Proof of theorem
Now, suppose this isn’t the case; for the claim to fail for this sequence it must be for a given $R>0$ that $M_{s_{i}}\cap B(0,R)$ has $m(s_{i})$ components for $s_{i}$ sufficiently negative. By the varifold convergence for $s_{i}$ sufficiently negative $m(s_{i})=m$. Using that the flow is smooth arguing exactly as in the lemma above where we showed the multiplicity of convergence was independent of sequence $s_{k}\to-\infty$ a compactness contradiction argument with the sheeting theorem gives that there will be no “new” sheets of $M_{s_{i}}\cap B(0,R)$ in that sheets of $M_{s_{i}}\cap B(0,R)$ correspond to components of $M_{s_{j}}\cap B(0,R)$ for $i,j$ large enough. Specifically, if this is not the case then because the flow is smooth there must be a sequence of times $t_{k}\to-\infty$ for which the flow isn’t locally sheeted over $M_{-\infty}$ (away from shrinking neighborhoods of a finite set of points in the sense described above), but one can extract a subsequence for which this must be true. With this in mind because $H_{t},H_{t}^{c}$ alternate across these sheets by the (strict) mean convexity of the flow we see that the multiplicity of convergence $m$ must be equal to two. The two connected components of $M_{s_{i}}\cap B(0,R)$ by set monotonicity would have to locally bound a slab in $H_{t}^{c}$, and because we assume strict mean convexity of the flow it must be the case $\partial H_{-\infty}$ lays strictly within. Now because that are no new sheets along $M_{s_{i}}$ for $i$ sufficiently negative we may sensically discuss them in between times $s_{i},s_{i+1}$, although it could be the case inbetween these times the are joined by bridges which may have propogated in from outside $B(0,R)$: we next rule these out. Supposing there are such bridges, one can see from the paragraph immediately below that because at the time $s_{i}$ there are no bridges in $B(0,R)$ by the set monotoncity of the flow they could only be given by necks which either bound solid cylindrical segments in $H_{t}$. Because $H_{t}$ is an increasing domain backwards in time at least some of the bridges must persist to time $s_{i+1}$, and as discussed more in the paragraph below such bridges cannot have moved far by set monotonicity of the flow. Considering $R$ slightly larger would then give a contradiction to the pressumption the claim fails, giving that $M_{-\infty}\cap B(0,R)\subset H_{t}^{c}$ for all sufficiently negative times. Because $M_{-\infty}$ is embedded so disconnects ${\mathbb{R}}^{3}$ and $m=2$ this gives for every $R>>0$ there is $s$ sufficiently negative so that $M_{t}\cap B(0,R)$ is disconnected for $t<s$. Because $H_{t}^{c}$ is an expanding domain under the flow (considered forward in time) then this gives that actually $M_{t}$ must be disconnected, giving a contradiction because we suppose $M_{t}$ is connected.
@@
So, for $R$ sufficiently large there must be a sequence as indicated above, where within $B(0,R)$ one can find bridges at least at the times $s_{i}$. Because in lemma 3.4 $\widetilde{M}\simeq S^{2}$ we see that $M_{s_{i}}\cap B(p_{j}(s_{i}),r_{j}(s_{i}))$ are homeomoprhic to a union of discs with a number of discs connected by cylindrical segments (justifying the use of the term “bridges”). These annuli could have other annuli nested within them, in the obvious sense, or bound solid cylinderical segments in $H_{t}$ or $H_{t}^{c}$ within the balls $B(p_{j}(s_{i}),r_{j}(s_{i}))$. Of course, the convergence to $M_{-\infty}$ is also in Hausdorff distance in bounded domains so as one looks back in time the curvature along these necks must blow up. In particular with the expository diversion in the last proof in mind if we had supposed a uniform curvature bound these would be ruled out immediately, going far enough back in time. For a bridge found at time $s_{i}$ as long it joins two sheets we have that for $s_{k}<s_{i}$ (indeed for all $t<s_{i}$) if the bridge laid in $B(p_{j}(s_{i}),r_{j}(s_{i}))$ it cannot be distance greater than $r_{j}(s_{i})$ from $p_{j}(s_{i})$, because otherwise a portion of $H_{t}^{c}$ would have to overlap with $H_{s_{i}}$ for some $t<s_{i}$ which is impossible considering the domain $H_{t}$ is increasing backwards under the flow; this is a coarse upper bound but suffices for our purposes as it keeps bridges we may find from sliding to spatial infinity farther back in time. Now apriori there is the possibility of a number of the bridges potentially disappearing as we follow them back in time corresponding to a drop in the number of sheets in the convergence. For instance those bounding solid cylinders segments in $H_{t}^{c}$ might bound on one side a ball in $H_{t}^{c}$ which may collapse considering the flow further back in time because this is a contracting domain, but if the neck persists it cannot move far. Because we have ruled out the “no bridges” case above and they can’t travel very far, it is the case that at least some of them do last as $t\to-\infty$ in $B(0,R)$ and without loss of generality we focus our attention on these “persistent” bridges.
@@
With this in mind, to start we can see that none of these cylindrical segments may bound a solid cylindrical segment in $H_{t}$. Suppose on the contrary one does, say contained in a ball $B(p_{j}(s_{i}),r_{j}(s_{i}))\subset B(0,R)$ where $R$ is sufficiently large. First note $\sup|A|^{2}\to\infty$ in $M_{s_{i}}\cap B(p_{j}(s_{i}),r_{j}(s_{i}))$ as $s_{i}\to\infty$, in particular it must do so along these connected cylindrical segments. Considering the flow is mean convex and so into the domain $H_{t}$, and the bridges contained in $B(p_{j}(s_{i}),r_{j}(s_{i}))$ are all contained in some large ball, say still $B(0,R)$ after adjusting $R$, even farther back in time implies that $\sup|A|^{2}=\infty$ within $M_{s_{i}}\cap B(p_{j}(s_{i}),r_{j}(s_{i}))\subset B(0,R)$, contradicting the smoothness of the flow. Alternately, perhaps one could argue using that since these solid cylinderical segments in $H_{t}$ should be expanding under the flow going back in time by the monotonicity of this set we would reach a contradiction to $M_{t}$ converging to $M_{-\infty}$. Similarly, we can rule out the appearance of any persistent annuli which have others nested within in $M_{s_{i}}\cap B(p_{j}(s_{i}),r_{j}(s_{i}))$.
As a result the only remaining cylindrical segments are those which bound a solid cylindrical segment in $H_{t}^{c}$. One can see (considering figure 1 below, where a similar idea is used in an argument below) then that consequently in any large ball $B(0,R)$ for $s_{i}$ sufficiently neagtive $M_{s_{i}}$ is the disjoint union of surfaces given by pairs of sheets bounding regions of $H_{t}$, with some pairs connected by such segments (so locally the regions are roughly speaking potentially punctured slabs). Our first claim in this case is that at most one of these pairs may persist as $s_{i}\to-\infty$, so the multiplicity of convergence is no more than 2. As in the no bridges case there will be no “new” pairs along the sequence $s_{i}$ for $i$ sufficiently large. Restricting ourselves to such $i$ because $M_{t}$ is connected for any two adjacent pairs there is a curve $\gamma$ in $M_{s_{i}}$ joining them, which should necessarily go outside $B(0,R)$. Now if we were not able to pick (in a continuous way) $\gamma_{t}$, where $\gamma_{0}=\gamma$ so that it remained in a fixed bounded region for more negative times it would be the case that for any $R>0$ if $i$ is large enough then $H_{s_{i}}\cap B(0,R)$ is disconnected – as in the no bridges case with the configuration of $M_{s_{i}}\cap B(0,R)$ in mind this would give that $M_{t}$ is disconnected giving a contradiction. So we may define $\gamma_{t}$ so it remains in a bounded region for more negative times, say within $B(0,\widetilde{R})$. Considering that $M_{s_{i}}$ eventually is sheeted over $M_{-\infty}$ in $B(0,\widetilde{R})$ as well $\gamma$ must go through a bridge far enough back in time, and such a neck by the choice of $\gamma_{t}$, as it goes between two “slabs” of $H_{t}$, must go along (one or a number of) of a nested necks or one which bounds a solid cylinder in $H_{t}$, particularly at least one of which must persist which was ruled out.
@@
Hence, for $i$ sufficiently large there must be some $j$ for which $H_{s_{i}}\cap B(p_{j}(s_{i}),r_{j}(s_{i}))\subset B(0,R)$ is a flattened solid torus laying in a slab. Again, this bridge cannot slide to infinity tracing it back farther in time, and will always be contained in a large ball (which again we may suppose will be $B(0,R)$). Note if $r_{j}$ is sufficiently small depending $M_{-\infty}$ and taking $s$ more negative if necessary, the width of the slab can be taken to be as small as one wishes – of course it is the case that the radii of all the balls where bridges occur in some fixed $B(0,R)$ tend to zero from earlier in the argument. Considering that the backwards limits $M_{t}$ and $M_{t+\delta}$ agree as $t\to-\infty$, as discussed in the argument for the previous lemma, then White’s expanding hole theorem [27] says that the cylindrical bridges cannot collapse in the backwards limit, giving a contradiction to $M_{t}$ varifold converging to $M_{-\infty}$. Alternately one can gain a contradiction using the mean convexity of the flow considering the sheets away from the bridges, because the domain $H_{t}$ increases backwards in time and so one can see at most one sheet could move towards $M_{-\infty}$ as $t$ is further decreased.
The upshot then is for any sequence $s_{i}\to-\infty$, a subsequence of $M_{s_{i}}$ must converge with multiplicity one to $M_{-\infty}$. Because the varifold convergence of $M_{t}\to M_{-\infty}$ is full this then implies the statement. $\square$
@@
By the mean convexity assumption the domain $H_{t}^{c}$ is decreasing in terms of set inclusion as we consider $t$ becoming more and more negative. We note that by using planes as barriers within the solid cylinders of lemma 3.3 that the number of ends of $M_{t}$ cannot decrease in finite time as we consider $t\to-\infty$ (if there were such a time, run the flow forward and use collections of planes as barriers to gain a contradiction), although as the bowl soliton shows this may occur in the limit (and the backwards limit can even be empty). If $M_{t}$ has two or more ends this issue doesn’t occur, however. If $M_{t}$ has two or more ends, we may consider properly embedded lines $\gamma(s,t)\subset M_{t}$ between any two of them, in that as $s\to-\infty$ $\gamma(s,t)$ lays on one end and as $s\to\infty$ $\gamma(s,t)$ lays on another. As $M_{t}$ is smooth such curves are defined for all $t$, and by lemma 3.3 and the mean convexity of $M_{t}$ any such choice of embedded line between two different ends must partially intersect a fixed compact region for all times $t$ giving that $M_{-\infty}$ is nonempty. Applying a similar argument within each of the solid half cylinders from lemma 3.3 one can also see that $M_{-\infty}$ will also have at least as many ends as $M_{t}$ does, although it could concievably have more.
Continuing to suppose $M_{t}$ has two or more ends, because $M_{-\infty}$ is a minimal surface of finite total curvature the ends of it are parallel by [17] (see also [21]), by which we mean $M_{-\infty}$ is ambiently isotopic to connect sums of parallel planes perhaps along with some compact surface: see figure 1 (also, figure 1 in [21]). This then implies that the ends of $M_{t}$ are parallel in the same sense. One can then readily see that if the number of ends is strictly greater than 2 then $\pi_{1}(H_{t}^{c})$ is not trivial, arguing as in lemma 3.7, giving a contradiction. $\square$
@@
(Alternate to lemma above) We point out that the statement above is actually stronger than we need to prove the theorem; it suffices to know that in the case that $M_{t}$ has two or more ends then $M_{-\infty}$ is nonempty and has a connected component with at least two ends. To see this, suppose that $M_{t}$ has two or more ends. Then from lemma 3.4 $\pi(H_{t})$ must be nontrivial. Considering a homotopically nontrivial curve in $\gamma\subset H_{0}$, because going backwards in time $H_{t}$ is an expanding domain it makes sense to consider $\gamma\subset H_{t}$ for $t<0$. If $M_{-\infty}$ is either empty or only contains components with one end which must be disjoint planes (since $M_{-\infty}$ has genus zero and finite total curvature) then $\gamma$ bounds a disc $D\subset H_{-\infty}$. After a potential slight perturbation of $D$ away from $M_{-\infty}$ then by the smooth convergence of $M_{t}\to M_{-\infty}$ for possibly very negative but finite $t$ $D\subset H_{t}$, giving a contradiction since the flow is smooth so $\gamma$ should stay homotopically nontrivial. Note the main result does imply the lemma above, aposteori.
@@
If the number of ends of $M_{t}$ is two, then from the argument of lemma 3.7 above $M_{-\infty}$ is nonempty and has at least two ends. By the halfspace theorem [14] either $M_{-\infty}$ is a single nonflat minimal surface or consists of a disjoint union of flat planes. Because $M_{-\infty}$ has finite total curvature it is homeomorphic to a genus $g$ surface with $k$ punctures for its $k$ ends – note that the number of ends could imaginably increase in the limit so that $k>2$, otherwise we would be done by, for instance, [23]. In fact, as a side remark we point out it can be written in a very concrete way as a compact surface of genus $g$ connected to a sum of parallel planes, up to isotopy (see figure 1 in [21]). Now, because the convergence of $M_{t}\to M_{-\infty}$ is smooth in compact domains by lemmas 3.5 and 3.6 we immediately get a contradiction if $g\geq 1$ to lemma 3.4. Since $g=0$ $M_{-\infty}$ is a punctured sphere so has genus 0 and, by the theorem of Lopez and Ros [19], must be a catenoid.
In the first, nonflat, case above then by lemma 3.2 $M_{t}$ must flow out of the catenoid in the manner indicated in the introduction, so we next rule out the latter case arguing similar to in the remark above. In the second case we see that $H_{-\infty}$ contains no homotopically nontrivial curves. On the other hand, by the set monotonicity of the flow $H_{t}\subset H_{-\infty}$ for all $t\in{\mathbb{R}}$, and in our case by lemma 3.7 $H_{t}$ does have such curves for all $t$; denote such a curve by $\gamma\subset H_{0}$ and by $D$ the immersed disc contained in $H_{-\infty}$ it bounds. By lemmas 3.5 and 3.6 the convergence of $H_{t}\to H_{-\infty}$ is smooth in compact domains, which implies that $D$, up to a slight perturbation, is contained in $H_{t}$ for $t$ sufficiently negative as well, giving a contradiction because $\gamma$ must remain homotopically nontrivial in $H_{t}$ since the flow is smooth. Alternately the proof of lemma 3.7 implies at least two ends of $M_{-\infty}$ must be connected, ruling out this case.

### 14 Mar 2026 v2

@@ 3. Proof of theorem
Now, suppose this isn’t the case; for the claim to fail for this sequence it must be for a given $R>0$ that $M_{s_{i}}\cap B(0,R)$ has $m(s_{i})$ components for $s_{i}$ sufficiently negative. By the varifold convergence for $s_{i}$ sufficiently negative $m(s_{i})=m$. Using that the flow is smooth arguing exactly as in the lemma above where we showed the multiplicity of convergence was independent of sequence $s_{k}\to$-$\infty$ a compactness contradiction argument with the sheeting theorem gives that there will be no “new” sheets of $M_{s_{i}}\cap B(0,R)$ in that sheets of $M_{s_{i}}\cap B(0,R)$ correspond to components of $M_{s_{j}}\cap B(0,R)$ for $i,j$ large enough. Specifically, if this is not the case then because the flow is smooth there must be a sequence of times $t_{k}\to$-$\infty$ for which the flow isn’t locally sheeted over $M_{-\infty}$ (away from shrinking neighborhoods of a finite set of points in the sense described above), but one can extract a subsequence for which this must be true. With this in mind because $H_{t},H_{t}^{c}$ alternate across these sheets by the (strict) mean convexity of the flow we see that the multiplicity of convergence $m$ must be equal to two. The two connected components of $M_{s_{i}}\cap B(0,R)$ by set monotonicity would have to locally bound a slab in $H_{t}^{c}$, and because we assume strict mean convexity of the flow it must be the case $\partial H_{-\infty}$ lays strictly within. Now because that are no new sheets along $M_{s_{i}}$ for $i$ sufficiently negative we may sensically discuss them in between times $s_{i},s_{i+1}$, although it could be the case in between these times they are joined by bridges which may have propagated in from outside $B(0,R)$: we next rule these out. Supposing there are such bridges, one can see from the paragraph immediately below that because at the time $s_{i}$ there are no bridges in $B(0,R)$ by the set monotonicity of the flow they could only be given by necks which either bound solid cylindrical segments in $H_{t}$. Because $H_{t}$ is an increasing domain backwards in time at least some of the bridges must persist to time $s_{i+1}$, and as discussed more in the paragraph below such bridges cannot have moved far by set monotonicity of the flow. Considering $R$ slightly larger would then give a contradiction to the presumption the claim fails, giving that $M_{-\infty}\cap B(0,R)\subset H_{t}^{c}$ for all sufficiently negative times. Because $M_{-\infty}$ is embedded so disconnects ${\mathbb{R}}^{3}$ and $m=2$ this gives for every $R>>0$ there is $s$ sufficiently negative so that $M_{t}\cap B(0,R)$ is disconnected for $t<s$. Because $H_{t}^{c}$ is an expanding domain under the flow (considered forward in time) then this gives that actually $M_{t}$ must be disconnected, giving a contradiction because we suppose $M_{t}$ is connected. (end of subclaim 1)
@@
So, for $R$ sufficiently large there must be a sequence as indicated above, where within $B(0,R)$ one can find bridges at least at the times $s_{i}$. Because in lemma 3.4 $\widetilde{M}\simeq S^{2}$ we see that $M_{s_{i}}\cap B(p_{j}(s_{i}),r_{j}(s_{i}))$ are homeomorphic to a union of discs with a number of discs connected by cylindrical segments (justifying the use of the term “bridges”). These annuli could have other annuli nested within them, in the obvious sense, or bound solid cylindrical segments in $H_{t}$ or $H_{t}^{c}$ within the balls $B(p_{j}(s_{i}),r_{j}(s_{i}))$. Of course, the convergence to $M_{-\infty}$ is also in Hausdorff distance in bounded domains so as one looks back in time the curvature along these necks must blow up. In particular with the expository lemma, lemma 3.6, in mind if we had supposed a uniform curvature bound these would be ruled out immediately, going far enough back in time.
Subclaim 2: WLOG, bridges are "persistent" and don’t wander far. For a bridge found at time $s_{i}$ as long it joins two sheets we have that for $s_{k}<s_{i}$ (indeed for all $t<s_{i}$) if the bridge laid in $B(p_{j}(s_{i}),r_{j}(s_{i}))$ it cannot be distance greater than $r_{j}(s_{i})$ from $p_{j}(s_{i})$, because otherwise a portion of $H_{t}^{c}$ would have to overlap with $H_{s_{i}}$ for some $t<s_{i}$ which is impossible considering the domain $H_{t}$ is increasing backwards under the flow; this is a coarse upper bound but suffices for our purposes as it keeps bridges we may find from sliding to spatial infinity farther back in time. Now a priori there is the possibility of a number of the bridges potentially disappearing as we follow them back in time corresponding to a drop in the number of sheets in the convergence. For instance those bounding solid cylinders segments in $H_{t}^{c}$ might bound on one side a ball in $H_{t}^{c}$ which may collapse considering the flow further back in time because this is a contracting domain, but if the neck persists it cannot move far. Because we have ruled out the “no bridges” case above and they can’t travel very far, it is the case that at least some of them do last as $t\to$-$\infty$ in $B(0,R)$ and without loss of generality we focus our attention on these “persistent” bridges, of which there must be at least one because we are supposing the convergence backwards in time is at least two. (end of subclaim 2)
@@
Subclaim 3: there are no solid or nested bridges. With subclaim 2 in hand, to start we can see that none of these cylindrical segments may bound a solid cylindrical segment in $H_{t}$. Suppose on the contrary one does, say contained in a ball $B(p_{j}(s_{i}),r_{j}(s_{i}))\subset B(0,R)$ where $R$ is sufficiently large. First note $\sup|A|^{2}\to\infty$ in $M_{s_{i}}\cap B(p_{j}(s_{i}),r_{j}(s_{i}))$ as $s_{i}\to$-$\infty$, in particular it must do so along these connected cylindrical segments. Considering the flow is mean convex and so into the domain $H_{t}$, and the bridges contained in $B(p_{j}(s_{i}),r_{j}(s_{i}))$ are all contained in some large ball, say still $B(0,R)$ after adjusting $R$, even farther back in time implies that $\sup|A|^{2}=\infty$ within $M_{s_{i}}\cap B(p_{j}(s_{i}),r_{j}(s_{i}))\subset B(0,R)$, contradicting the smoothness of the flow. Alternately, perhaps one could argue using that since these solid cylindrical segments in $H_{t}$ should be expanding under the flow going back in time by the monotonicity of this set we would reach a contradiction to $M_{t}$ converging to $M_{-\infty}$. Similarly, we can rule out the appearance of any persistent annuli which have others nested within in $M_{s_{i}}\cap B(p_{j}(s_{i}),r_{j}(s_{i}))$. (end of subclaim 3)
[…]
@@
Hence, for $i$ sufficiently large there must be some $j$ for which $H_{s_{i}}\cap B(p_{j}(s_{i}),r_{j}(s_{i}))\subset B(0,R)$ is a flattened punctured ball laying in a slab. Again, this bridge cannot slide to infinity tracing it back farther in time, and will always be contained in a large ball (which again we may suppose will be $B(0,R)$). Note if $r_{j}$ is sufficiently small depending $M_{-\infty}$ and taking $s$ more negative if necessary, the width of the slab can be taken to be as small as one wishes – of course it is the case that the radii of all the balls where bridges occur in some fixed $B(0,R)$ tend to zero from earlier in the argument. Considering that the backwards limits $M_{t}$ and $M_{t+\delta}$ agree as $t\to$-$\infty$, as discussed in the argument for the previous lemma, then White’s expanding hole theorem [28] says that the cylindrical bridges cannot collapse in the backwards limit, giving a contradiction to $M_{t}$ varifold converging to $M_{-\infty}$. Alternately one can gain a contradiction using the mean convexity of the flow considering the sheets away from the bridges, because the domain $H_{t}$ increases backwards in time and so one can see at most one sheet could move towards $M_{-\infty}$ as $t$ is further decreased. (end of subclaim 4)
To summarize we showed that if the number of sheets in the backwards convergence was with multiplicity greater than one, then there must be at least one persistent bridge joining them using the set monotonicity of the flow. Again using the set monotonicity of the flow though we could rule out such bridges, implying that for any sequence $s_{i}\to$-$\infty$, a subsequence of $M_{s_{i}}$ must converge with multiplicity one, and hence smoothly, to $M_{-\infty}$. Because the varifold convergence of $M_{t}\to M_{-\infty}$ is full this then implies the statement. $\square$
@@
By the mean convexity assumption we observe the domain $H_{t}^{c}$ is decreasing in terms of set inclusion as we consider $t$ becoming more and more negative, or in other words $H_{t}^{c}\subset H_{0}^{c}$ for all $t$ negative. We note that by using planes as barriers within the solid cylinders of lemma 3.3 that the number of ends of $M_{t}$ cannot decrease in finite time as we consider $t\to$-$\infty$ (if there were such a time, run the flow forward and use collections of planes intersecting the solid half cylinders as barriers to gain a contradiction – no noncompact maximum principles are required as the cross sections of solid cylinders are compact), although as the bowl soliton shows this may occur in the limit and in fact the backwards limit can even be empty. If $M_{t}$ has two or more ends this issue doesn’t occur, however. If $M_{t}$ has two or more ends, we may consider properly embedded lines $\gamma(s,t)\subset M_{t}$ between any two of them, in that as $s\to$-$\infty$ $\gamma(s,t)$ lays on one end and as $s\to\infty$ $\gamma(s,t)$ lays on another. As $M_{t}$ is smooth such curves are defined for all $t$, and by lemma 3.3 and the mean convexity of $M_{t}$ any such choice of embedded line between two different ends must intersect $H_{0}^{c}\cap B(0,R)$, $R>>0$, for all times $t<0$ giving that $M_{-\infty}$ is nonempty. Applying a similar argument within each of the solid half cylinders from lemma 3.3 one can also see that $M_{-\infty}$ will also have at least as many ends as $M_{t}$ does, although it could conceivably have more.
Continuing to suppose $M_{t}$ has two or more ends, because $M_{-\infty}$ is a minimal surface of finite total curvature the ends of it are parallel by [17] (see also [21]), by which we mean $M_{-\infty}$ is ambiently isotopic to connect sums of parallel planes perhaps along with some compact surface: see figure 1 (also, figure 1 in [21]). This then implies that the ends of $M_{t}$ are parallel in the same sense. One can then readily see that if the number of ends is strictly greater than 2 then $\pi_{1}(H_{t}^{c})$ is not trivial, giving a contradiction to lemma 3.4. $\square$
@@
(Alternate to lemma above) We point out that the statement above is actually stronger than we need to prove the theorem; it suffices to know that in the case that $M_{t}$ has two or more ends then $M_{-\infty}$ is nonempty and has a connected component with at least two ends. To see this, suppose that $M_{t}$ has two or more ends. Then from lemma 3.4 $\pi_{1}(H_{t})$ must be nontrivial. Considering a homotopically nontrivial curve in $\gamma\subset H_{0}$, because going backwards in time $H_{t}$ is an expanding domain it makes sense to consider $\gamma\subset H_{t}$ for $t<0$. If $M_{-\infty}$ is either empty or only contains components with one end which must be disjoint planes (since $M_{-\infty}$ has genus zero and finite total curvature) then $\gamma$ bounds a disc $D\subset H_{-\infty}$. After a potential slight perturbation of $D$ away from $M_{-\infty}$ then by the local smooth convergence of $M_{t}\to M_{-\infty}$ for possibly very negative but finite $t$ $D\subset H_{t}$, giving a contradiction since the flow is smooth so $\gamma$ should stay homotopically nontrivial. Note the main result does imply the lemma above, a posteori.
@@
If the number of ends of $M_{t}$ is two, then from the argument of lemma 3.9 above $M_{-\infty}$ is nonempty and has at least two ends. By the halfspace theorem [14] either $M_{-\infty}$ is a single nonflat minimal surface or consists of a disjoint union of flat planes. Because $M_{-\infty}$ has finite total curvature it is homeomorphic to a genus $g$ surface with $k$ punctures for its $k$ ends – note that the number of ends could imaginably increase in the limit so that $k>2$, otherwise we would be done by, for instance, [23]. In fact, as a side remark we point out it can be written in a very concrete way as a compact surface of genus $g$ connected to a sum of parallel planes, up to isotopy (see figure 1 in [21]). Now, because the convergence of $M_{t}\to M_{-\infty}$ is smooth in compact domains by lemmas 3.7 and 3.8 we immediately get a contradiction if $g\geq 1$ to lemma 3.4. Since $g=0$ $M_{-\infty}$ is a punctured sphere so has genus 0 and, by the theorem of Lopez and Ros [19], must be a catenoid.
In the first, nonflat, case above then by lemma 3.2 $M_{t}$ must flow out of the catenoid in the manner indicated in the introduction, so we next rule out the latter case arguing similar to in the remark above. In the second case we see that $H_{-\infty}$ contains no homotopically nontrivial curves. On the other hand, by the set monotonicity of the flow $H_{t}\subset H_{-\infty}$ for all $t\in{\mathbb{R}}$, and in our case by lemma 3.9 $H_{t}$ does have such curves for all $t$; denote such a curve by $\gamma\subset H_{0}$ and by $D$ the immersed disc contained in $H_{-\infty}$ it bounds. By lemmas 3.7 and 3.8 the convergence of $H_{t}\to H_{-\infty}$ is smooth in compact domains, which implies that $D$, up to a slight perturbation, is contained in $H_{t}$ for $t$ sufficiently negative as well, giving a contradiction because $\gamma$ must remain homotopically nontrivial in $H_{t}$ since the flow is smooth. Alternately the proof of lemma 3.9 implies at least two ends of $M_{-\infty}$ must be connected, ruling out this case.